for the following function (5 Marks) dy 5. (b) Use implicit differentiation to find dx 3x* - 6 ln(2xy) = 7-5e24-sy dy
1. (b) Evaluate: lim r? + 4x -12 (3-12x +1 (5 Marks)
Answer #1

(a)3x* – 61n (2x*y)= 7 – 5e2x=sy Differentiating with respect x, we get 3(4x) - 6 2x +6Xy] =0–s[2010-> 562x-Sy dy 12x2 2x-5 y